OpenStudy (anonymous):

Find the equation in slope-intercept form of the line that passes through the points (-2,4) and (6, 2). A) y = 4x + 12 B) y = 4x - 18 C) y =1/4x +9/2 D) y = -1/4x +7/2

OpenStudy (anonymous):

Do you know which one is x1 and x2?

OpenStudy (anonymous):

Because I forget which one is which

OpenStudy (anonymous):

x1 issssss the 2nd one of the first one i think

OpenStudy (anonymous):

x1,y1 x2,y2

OpenStudy (anonymous):

ohhhhh so that's how you do this hahaha

OpenStudy (anonymous):

Still need me to help? :P

OpenStudy (anonymous):

yeah i got 2 over negative 8 and now i dunno what to do :|

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

Okay i'll show you how to do this: y = mx + b So since you know which points are x1,y1 & x2,y2 you'd do M = y2-y1 ------- x2-x1 The y-intercept b of the line is the value of y at the point where the line crosses the y axis. Since for point (x1, y1) we have y1 = mx1 + b, the y-intercept b can be calculated by: b = y1 - mx1 But the answer is C.

OpenStudy (anonymous):

dang it lol

OpenStudy (anonymous):

You were closer than a and b :P